### v1

# ID-Pose: Sparse-view Camera Pose Estimation by Inverting Diffusion Models

Abstract
Given sparse views of an object, estimating their camera poses is a long-standing and intractable problem. We harness the pre-trained diffusion model of novel views conditioned on viewpoints (Zero-1-to-3). We present ID-Pose which inverses the denoising diffusion process to estimate the relative pose given two input images. ID-Pose adds a noise on one image, and predicts the noise conditioned on the other image and a decision variable for the pose. The prediction error is used as the objective to find the optimal pose with the gradient descent method. ID-Pose can handle more than two images and estimate each of the poses with multiple image pairs from triangular relationships. ID-Pose requires no training and generalizes to real-world images. We conduct experiments using high-quality real-scanned 3D objects, where ID-Pose significantly outperforms state-of-the-art methods. https://xt4d.github.io/id-pose/.
Figure 1: ID-Pose estimates the relative camera pose of two input images. It uses one image as a reference and the other image as the target, where the target image is added with Gaussian noise. Assume there is a relative transformation from the camera pose of the reference to the target. ID-Pose utilizes the noise predictor from Zero-1-to-3 [8] to compute the noise given the noisy target image, the reference image, and the relative pose. The prediction error is used as the objective of the relative pose. ID-Pose then applies the gradient descent method to find the optimal relative pose that connects the two views.
1 Introduction
Estimating camera poses of images is an important task for 3D understanding, reconstruction [11, 6, 24], and generation [21]. Given sparse views of an object, it is often intractable to estimate their poses. Traditional methods such as Structural-from-Motion (SfM) [17] require dense local features [9, 19, 1] that are unavailable with sparsely sampled images. Recent methods train neural networks to directly predict poses given images [10, 22, 18, 7]. However, the generalization ability of these methods is limited by the scale of training data, which requires expensive dense image collection [15] or exhaustive pose annotating [5].
We solve the camera pose estimation problem by harnessing the pre-trained diffusion model of novel views conditioned on viewpoints (Zero-1-to-3) [8]. We present ID-Pose which inverses the denoising diffusion process to estimate the relative pose of two input images. ID-Pose leverages the pre-trained diffusion model, and maintains the relative pose as the optimization variable. It is conditioned on one image and the pose to predict the noise added to the other image. The error of the prediction is used as the objective to update the pose with the gradient descent method. To prevent local minimum, we propose an initialization strategy that first samples a number of candidature poses with a few steps of optimization and then selects the one with the minimum prediction error of noise. We further extend ID-Pose to effectively find relative poses of more than two images. The idea is to estimate a pose with multiple image pairs. Within a group of three images, we can use the poses of two image pairs to present the pose of the third pair. Therefore, the prediction error of the third pair can be used to update the two poses alternately. ID-Pose is a zero-shot method that generalizes to open-world images by leveraging the diffusion model pre-trained on large-scale data [8, 16]. We conduct experiments on high-quality real-scanned 3D objects, where ID-Pose significantly outperforms state-of-the-art methods.
2 Related Work
Camera Pose Estimation. Estimating camera poses of images is a long-standing problem. Traditional methods extract local features [9, 19, 1] of input images and find their matches to estimate poses. When images are sparsely captured, they often fail to estimate poses due to insufficient feature matches. Recent learning based methods [10, 22, 18, 7] use neural networks to predict the relative poses of image pairs. These methods require a large number of pose annotations, which are usually estimated with dense views of an object [15]. As training data is expensive to obtain, the neural networks are limited to benchmarks and hard to handle out-of-distribution images.
Image Diffusion Models. Diffusion models [4] have been widely used for image generation tasks. They learn a neural network to denoise an image, that takes a noisy image to predict the noise values mixed in the image. As the noise is random, the model essentially predicts the image and subtracts it from the noisy image to obtain the noise values. By progressive denoising from pure noise, a clear image can be generated. Latent diffusion models [16] denoise latent representations of images, which reduce the model complexity and improve the efficiency and stability. Recent models are able to condition on other inputs to control the generation, for example, natural language and sketches [23, 12]. Stable Diffusion [16] is a latent diffusion model that generates high-quality images based on input texts by training on over a hundred million image-text pairs. A view-conditioned diffusion model, Zero-1-to-3 [8], uses one input image of an object and a relative pose to generate the novel view of the object under the new viewpoint. It is fine-tuned from Stable Diffusion with pose-annotated images rendered using Objaverse dataset [3]. It achieves prominent generation results with real-world images. We consider to inverse the denoising process of Zero-1-to-3, that uses two images of the same object to predict their relative camera pose. This enables zero-shot pose estimation given a sparse set of images.
3 Methodology
Our task is to estimate camera poses of images that are sparsely captured around an object. We present ID-Pose, a zero-shot method that takes two images as input and finds their relative camera pose. The idea is to invert the denoising process of the pre-trained novel view generation model, i.e., Zero-1-to-3 [8]. We extend ID-Pose to handle more than two images by inverting each of the poses with multiple image pairs. In this section, we first introduce the background knowledge of diffusion models, and then describe the details of the ID-Pose method.
3.1 Preliminary
Let $x$ be an image, and $z$ be the latent map of $x$. Given a Gaussian noise $\epsilon$, one can add a small portion of $\epsilon$ to $z$ step-by-step, and finally turn $z$ into $\epsilon$. At each step $t$, a noisy latent map $z(\epsilon,t)$ is obtained by mixing $z$ and $\epsilon$ with proportion scales controlled by $t$. Latent diffusion models [16] learn a step-wise noise predictor $f(z(\epsilon,t),t)$ that predicts the noise added to $z(\epsilon,t)$. To generate an image, a latent diffusion model starts from a noise $\epsilon$ and assumes there is a $z$ underlying the $\epsilon$. It then predicts and removes the noise step-by-step. After a fixed number of steps, it recovers the latent map $z$, which is then input to an image decoder to obtain the resulting image.
Stable Diffusion [16] is a latent diffusion model that generates images conditioned on text. Its noise predictor is therefore written as $f(z(\epsilon,t),text,t)$. The condition $text$ is mapped to CLIP [14] embedding tokens which are cross-attended with the hidden states of the predictor network.
Zero-1-to-3 [8] is fine-tuned from Stable Diffusion that generates novel views of an object conditioned on a reference view $x_{c}$ and a relative camera pose $p$. Its noise predictor is written as $f(z(\epsilon,t),x_{c},p,t)$. The reference view $x_{c}$ is mapped to image tokens with the CLIP image encoder. The camera pose $p$ is expressed in a spherical coordinate system where the position of a point is specified by three numbers: polar angle, azimuth angle, and radius. The three values of the pose $p$ are concatenated with the $x_{c}$ tokens as conditions to the predictor $f$. Zero-1-to-3 is fine-tuned with large-scale images rendered from 3D objects but generalizes to real-world images.
3.2 ID-Pose
Let $f(z(\epsilon),x,p)$ be the pre-trained noise predictor of Zero-1-to-3, where the diffusion step $t$ is omitted for simple notations. Given two images $x_{0}$ and $x_{1}$, we inverse the denoising process to estimate their relative pose, which is represented in spherical coordinates (polar, azimuth, radius). Let $x_{0}$ be the reference view and $x_{1}$ be the target view. We add a noise $\epsilon$ to $x_{1}$ and use $f$ to predict the noise. Assume $p$ is the pose transformation from $x_{0}$ to $x_{1}$, the predicted noise is defined as a function of $p$:
$\hat{\epsilon}(p\,;\,x_{0},x_{1})=f(z_{1}(\epsilon),x_{0},p),$ (1)
where $z_{1}(\epsilon)$ is the noisy latent map of $x_{1}$. We define the noise error of the prediction given $p$ as:
$l(p\,;\,x_{0},x_{1})=||\hat{\epsilon}(p\,;\,x_{0},x_{1})-\epsilon||_{2}^{2}.$ (2)
We also consider that $x_{1}$ acts as the reference view and $x_{0}$ acts as the target view. Let $p^{\prime}$ be the reversed transformation of $p$ (from $x_{1}$ to $x_{0}$). The predicted noise of $p^{\prime}$ is:
$\hat{\epsilon}(p^{\prime}\,;\,x_{1},x_{0})=f(z_{0}(\epsilon),x_{1},p^{\prime}),$ (3)
where $z_{0}(\epsilon)$ is the noisy latent map of $x_{0}$. The noise error of the prediction given $p^{\prime}$ is:
$l(p^{\prime}\,;\,x_{1},x_{0})=||\hat{\epsilon}(p^{\prime}\,;\,x_{1},x_{0})-\epsilon||_{2}^{2}.$ (4)
We randomly choose $l(p\,;\,x_{0},x_{1})$ or $l(p^{\prime}\,;\,x_{1},x_{0})$ as the objective function $g(p)$. The inversion process iteratively updates $p$ by the gradient descent method:
$p^{(k+1)}=p^{(k)}-\alpha\nabla g(p^{(k)})$ (5)
where $\alpha$ is the step factor, $k$ represents the current updating step, and $k+1$ represents the next step. The iteration stops when $p$ converges or $k$ exceeds a maximum step count. The intuition is that, $p$ reaches the optimum once the model can predict the noise with the minimum error.
The initialization of $p$ is crucial as optimization can be often stuck into local minimums. For a near-symmetric object, the opposite view of one view can have the highest similarity among the other views. Under this situation, if $p$ is initialized near the opposite view, it will be likely stuck into a local minimum. To alleviate this problem, we use a sampling approach to find reliable initialization. We uniformly sample a number of pose candidates on a sphere around the origin point. The relative radius of the candidature poses is set to zero. For each candidate, we update the pose for a few steps with Equation 5. We then evaluate those poses with the pairwise error:
$l_{0,1}(p)=l(p\,;\,x_{0},x_{1})+l(p^{\prime}\,;\,x_{1},x_{0}),$ (6)
which combines the two noise errors to measure the stability of $p$ connecting $x_{0}$ and $x_{1}$. From all the candidates, we pick the one with the minimum pairwise error as the initial pose and further update it until convergence.
3.3 ID-Pose with More Views
Given $n$ ($n>2$) images, a naive strategy is breaking them into $n-1$ consecutive pairs, and inverting the pose for each of them independently. However, there are many unused image pairs that can help to find these poses. We consequently propose to estimate each of the relative poses with multiple image pairs from triangular relationships. Suppose the input images are $x_{0}$, $x_{1}$, …, $x_{n-1}$, we maintain $n-1$ relative poses $p_{1}$, $p_{2}$, …, $p_{n-1}$ as variables, where $p_{i}$ represents the relative transformation from $x_{0}$ to $x_{i}$. Firstly, we describe the initialization process. For each $p_{i}$, we uniformly sample $m$ candidates and update them with a few steps. Instead of picking the best one, we keep those candidates as a group $\{p_{i,u}|u\in[0,m)\}$. Given two poses $p_{i}$ and $p_{j}$ where $i\neq j$, we define their triangular error as:
$l_{i,0,j}(p_{i},p_{j})=l_{0,i}(p_{i})+l_{0,j}(p_{j})+l_{i,j}(p^{\prime}_{i}+p_{j}),$ (7)
where $p^{\prime}_{i}+p_{j}$ is the relative transformation from $x_{i}$ to $x_{j}$. The triangular error sums up three pairwise errors, which indicates the instability of poses $p_{i}$ and $p_{j}$ to connect views $x_{0}$, $x_{i}$, $x_{j}$. Then, given a candidate $p_{i,u}$ of $p_{i}$, we iterate the candidates $p_{j,v}$ of $p_{j}$ and calculate triangular errors ($m$ times). We define the minimum error of $p_{i,u}$ pairing with $p_{j}$ as:
$e_{i,u,j}=\min_{v\in[0,m)}l_{i,0,j}(p_{i,u},p_{j,v}).$ (8)
We obtain $e_{i,u}=\sum_{j\neq i}e_{i,u,j}$ as the instability of $p_{i,u}$. We select $p_{i,u}$ with lowest instability as the initialization of $p_{i}$.
Next, we describe the inversion process. Suppose all poses are initialized, we randomly pick two indices $i$ and $j$ where $0\leq i\neq j<n$. We consider three situations. If $i=0$, we update $p_{j}$ with the noise error $l(p_{j}\,;\,x_{0},x_{j})$. If $j=0$, we update $p_{i}$ with the noise error $l(p^{\prime}_{i}\,;\,x_{i},x_{0})$. If both $i>0$ and $j>0$, we will have two poses $p_{i}$ and $p_{j}$ to update. We fix one and update the other with the noise error $l(p^{\prime}_{i}+p_{j}\,;\,x_{i},x_{j})$.
4 Experiments
4.1 Experimental Settings
Data. We conduct experiments on two datasets: OmniObject3D [20] and Amazon Berkeley Objects (ABO) [2]. These datasets are only used for testing, and none of the experimental methods are trained on them. OmniObject3D is a dataset of high-quality real-scanned 3D objects. It includes 190 categories with 6,000 scanned objects. We select 12 categories “house”, “chair”, “table”, “bed”, “ornaments”, “shoe”, “scissor”, “guitar”, “laundry detergent”, “dumbbell”, “dinosaur”, and “glasses”. We select the first 5 scanned objects for each category if available. The total number of objects is 54. For each object, we use all the Blender rendered images as a view pool, and change their background color to white. ABO is a dataset of 3D Amazon.com products. We select 10 representative objects: “bag”, “bar stool chair”, “ergonomic chair”, “sofa”, “couch”, “desk”, “shelf”, “bed”, “fan”, and “lamp”. For each object, we place its geometry center at the origin point. We then setup a camera with FOV=$60^{\circ}$ looking at the origin, and randomly place it on a sphere to render 30 images. The radius of the sphere is set to a value that the object can be fully captured with enough image margin. The background is set to white during rendering. The rendered images form a view pool of an object. The ground truth poses are known and associated with the images. Given a view pool, we randomly pick out $n$ images as a testing sample. We experiment with $n$ = $2,4,6,8$. For OmniObject3D, we repeat the picking 3 times. We make $648$ testing samples in total containing 2,592 relative poses to estimate. For ABO, we repeat the picking 10 times. We make $400$ testing samples in total containing 1,600 relative poses to estimate.
Baseline Settings. We use RelPose++ [7] as the baseline method, which is trained using Common Objects 3D (v2) dataset [15]. There are two released model checkpoints. One is trained using raw images, the other is trained using masked images with clean background. We evaluate both of them in the experiments.
ID-Pose Settings. We use the checkpoint of Zero-1-to-3 [8] trained with 105000 iterations. We preprocess input images to have white background and contain enough margin, which fit the input distribution of the noise predictor. Given an image, we use a segmentation mask to clear the background, and use a window to crop out the object at its center. The size of the window is initially set to the largest bounding box of segmentation masks of all input images. The size is then scaled 1.5x to ensure that there is enough margin in the cropped images. The hyper-parameters used in ID-Pose are shown in Table 1. We additionally test ID-Pose without using triangular relationships.
Initialization Stage Optimization Stage
Noise $\epsilon$ fixed random (each iteration)
Noise level 0.8 0.8
Step factor $\alpha$ 10.0 1.0
Batch size 16 1
Iterations 10 (each pose) $600\times(n-1)$
Table 1: The hyper-paramters of ID-Pose.
Metrics. Given $n$ images and the camera pose of one image. We estimate relative poses and obtain the poses for the other $n-1$ images. The poses are transformed into world coordinate system. Following RelPose++[7], we obtain relative rotations of the estimated poses and compute the angular differences to the ground truth relative rotations. We report the rotation accuracy which is the proportion of the angular differences less than 15 or 30 degree. We also compute the $l_{2}$ distances from the estimated camera positions to the ground truth positions. The distances are normalized by dividing the corresponding ground truth radius. We report the camera position accuracy which is the proportion of distances less than 20%.
4.2 Evaluation Results
$n=2$ $n=4$ $n=6$ $n=8$
Method $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA
RelPose++ w/o Mask 5.56 14.20 8.64 4.12 12.35 7.20 3.70 12.72 8.27 4.41 12.61 9.08
RelPose++ 6.17 11.11 12.96 4.32 11.73 12.76 3.70 11.85 11.60 4.50 11.90 11.46
ID-Pose w/o Tri 48.77 56.79 45.06 48.77 59.88 44.65 48.77 60.25 44.57 49.65 61.38 45.06
ID-Pose 48.77 56.79 45.06 49.79 61.11 46.50 50.12 63.33 46.42 52.38 65.26 48.06
Table 2: Results on the OmniObject3D testing samples. “$15^{\circ}$”: rotation accuracy by $15^{\circ}$. “$30^{\circ}$”: rotation accuracy by $30^{\circ}$. “PA”: camera position accuracy. The numbers are shown in 100%.
$n=2$ $n=4$ $n=6$ $n=8$
Method $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA
RelPose++ w/o Mask 2.00 6.00 2.00 2.00 7.33 1.67 3.00 7.40 3.20 2.86 7.86 3.43
RelPose++ 2.00 6.00 1.00 2.00 7.00 1.67 4.00 7.60 4.00 3.43 7.29 3.86
ID-Pose w/o Tri 76.00 80.00 73.00 75.00 78.33 74.00 77.60 80.00 75.60 77.71 80.57 75.86
ID-Pose 76.00 80.00 73.00 77.00 79.33 74.00 78.00 81.40 75.60 78.71 82.14 77.14
Table 3: Results on the ABO testing samples.
We report evaluation results aggregated by the number of input views. The results on OmniObject3D and ABO samples are shown in Table 2 and Table 3, respectively. ID-Pose methods significantly outperform RelPose++ for all different $n$, which demonstrate advanced generalization capability. RelPose++ methods report much lower metrics compared to the testing results on the CO3D dataset [7]. The reason might be that the models of RelPose++ overfit to CO3D data and fail to generalize to out-of-distribution images. With two input images ($n=2$), ID-Pose w/o Tri and ID-Pose report the same results as they behave identically. With more input images ($n>2$), ID-Pose outperforms ID-Pose w/o Tri as extra image pairs are used to estimate poses. By adding input images, the performance of ID-Pose increase as more image pairs are used.
We visualize examples of estimated camera poses in the world coordinate system. We compare ID-Pose and Relpose++ with ground truth. Assume that the true camera pose of the reference image is known. For RelPose++, we transform all predicted poses with a matrix that aligns the first prediction to the reference. In Figure 2, we input four views of the “house_001” from OmniObject3D, where the front view acts as the reference. Compared with the ground truth, ID-Pose finds all three relative poses with small errors. RelPose++ accurately predicts the relative poses of the two side views, but makes a large prediction error on the opposite view. In Figure 3, we input four views of the “ornaments_001” from OmniObject3D, where the “left-face” view acts as the reference. This is a hard case for both ID-Pose and RelPose++. ID-Pose finds the approximate poses but with large errors compared to the ground truth. RelPose++ predicts poses with even larger errors. In Figure 4, we input four views of an ABO chair object. ID-Pose accurately finds the camera poses, while RelPose++ fails to distinguish the differences between these images. This failure happens to almost all the ABO testing samples. RelPose++ is probably hard to handle out-of-distribution images. This exposes the weakness of the fitting based methods, that the generalization capability is limited by the amount of annotated training data. ID-Pose leverages the model pre-trained on large-scale images, which demonstrates better generalization ability.
Figure 2: Visualization of camera poses with a sample of views from OmniObject3D “house_001”.
Figure 3: Visualization of camera poses with a sample of views from OmniObject3D “ornaments_001”.
Figure 4: Visualization of camera poses with a sample of views from an ABO chair.
4.3 Qualitative Results on In-the-wild Images
We qualitatively show the results of ID-Pose on in-the-wild images. We find three objects in our daily workplace: a toy duck, a chair, and a foosball table. We use a smartphone to capture 4-6 images for each of them. We use the DIS method [13] to segment out the objects and clean the background of the images to white. We select $n$ ($n=2,3,4$) images as input views to estimate the relative poses. To visualize the results, we use the LumaAI APP [^1] https://lumalabs.ai/ to reconstruct 3D meshes of the objects. We then choose one image as the reference and manually find its camera pose using the 3D mesh. With the ground truth camera pose of the reference image, we transform the estimated poses of the target views into the world coordinate system. We show an example of the results of the chair in Figure 5. More results can be interactively visualized on the project website [^2] https://xt4d.github.io/id-pose/. The estimated poses cannot align the views to the 3D object perfectly. One reason is that the true cameras may not look at the same point in the world, and their “up” vectors can be varied. For in-the-wild images, pose estimation with the assumption of spherical coordinates will introduce additional errors. This is a weakness of ID-Pose.
Figure 5: Visualization of estimated camera poses of sparsely captured views around a chair. The mesh is reconstructed using LumaAI APP. The view in red acts as the reference view, where the camera pose is manually found using the mesh.
5 Conculsion & Limitations
In this paper, we propose ID-Pose which inverses the denoising diffusion process to estimate camera poses of sparse input views. ID-Pose only utilizes the pre-trained Zero-1-to-3 model [8] without additional training. Given two images, ID-Pose adds a noise on one image, and predicts the noise conditioned on the other image and a decision variable for the relative pose. The prediction error is used as the objective to find the optimal pose with the gradient descent method. ID-Pose can handle more than two images. It estimates each of the poses with additional pairs of images from triangular relationships. We conduct experiments using high-quality real-scanned 3D objects, where ID-Pose significantly outperforms state-of-the-art methods.
Limitations. The running time of ID-Pose is significantly higher than neural fitting based methods. It requires multiple steps of pose updating by inference through large diffusion networks. The average time for running 100 steps is about 30 seconds on an NVIDIA V100 GPU. This is time-consuming and the method needs efficiency improvement. Due to the limitation of Zero-1-to-3, ID-Pose assumes that all cameras look at the same point and have no rotation around optical direction. This will introduce estimation errors for in-the-wild images. To extend the capability, one possible solution is to fine-tune the novel view diffusion model with the input of 6DOF poses rather than spherical coordinates.

### v2

@@
 Abstract
-Given sparse views of an object, estimating their camera poses is a long-standing and intractable problem.
-We harness the pre-trained diffusion model of novel views conditioned on viewpoints (Zero-1-to-3).
+Given sparse views of a 3D object, estimating their camera poses is a long-standing and intractable problem.
+Toward this goal, we consider harnessing the pre-trained diffusion model of novel views conditioned on viewpoints (Zero-1-to-3).
 We present ID-Pose which inverses the denoising diffusion process to estimate the relative pose given two input images.
-ID-Pose adds a noise on one image, and predicts the noise conditioned on the other image and a decision variable for the pose.
-The prediction error is used as the objective to find the optimal pose with the gradient descent method.
-ID-Pose can handle more than two images and estimate each of the poses with multiple image pairs from triangular relationships.
-ID-Pose requires no training and generalizes to real-world images.
-We conduct experiments using high-quality real-scanned 3D objects, where ID-Pose significantly outperforms state-of-the-art methods.
-https://xt4d.github.io/id-pose/.
+ID-Pose adds a noise to one image, and predicts the noise conditioned on the other image and a hypothesis of the relative pose.
+The prediction error is used as the minimization objective to find the optimal pose with the gradient descent method.
+We extend ID-Pose to handle more than two images and estimate each pose with multiple image pairs from triangular relations.
+ID-Pose requires no training and generalizes to open-world images.
+We conduct extensive experiments using casually captured photos and rendered images with random viewpoints.
+The results demonstrate that ID-Pose significantly outperforms state-of-the-art methods.
+[Project Page: https://xt4d.github.io/id-pose-web/]
 Figure 1: ID-Pose estimates the relative camera pose of two input images.
-It uses one image as a reference and the other image as the target, where the target image is added with Gaussian noise.
+It uses one image as the reference and the other image as the target, where the target image is added with Gaussian noise.
 Assume there is a relative transformation from the camera pose of the reference to the target.
-ID-Pose utilizes the noise predictor from Zero-1-to-3 [8] to compute the noise given the noisy target image, the reference image, and the relative pose.
-The prediction error is used as the objective of the relative pose.
+ID-Pose utilizes the noise predictor of Zero-1-to-3 [13] to compute the noise given the noisy target image, the reference image, and the relative pose.
+The prediction error is used as the minimization objective regarding the relative pose.
 ID-Pose then applies the gradient descent method to find the optimal relative pose that connects the two views.
 1 Introduction
-Estimating camera poses of images is an important task for 3D understanding, reconstruction [11, 6, 24], and generation [21].
-Given sparse views of an object, it is often intractable to estimate their poses.
-Traditional methods such as Structural-from-Motion (SfM) [17] require dense local features [9, 19, 1] that are unavailable with sparsely sampled images.
-Recent methods train neural networks to directly predict poses given images [10, 22, 18, 7].
-However, the generalization ability of these methods is limited by the scale of training data, which requires expensive dense image collection [15] or exhaustive pose annotating [5].
-We solve the camera pose estimation problem by harnessing the pre-trained diffusion model of novel views conditioned on viewpoints (Zero-1-to-3) [8].
+Estimating camera poses of images that depict a 3D object is important to shape understanding [29, 17], reconstruction [16, 34, 8, 37], and generation [23, 9, 32, 12].
+With densely scanned images, this problem is well-addressed with Structural-from-Motion (SfM) [26].
+However, given casually captured images, which are often sparse in terms of viewpoints, SfM methods frequently encounter convergence difficulties due to the insufficient overlapping appearances between different views.
+With a growing interest in unveiling 3D from daily images, addressing the challenge of estimating camera poses in sparse-view settings becomes an increasingly urgent task.
+Recent methods train feed-forward neural networks to predict camera poses given images [15, 35, 27, 10].
+These methods require training with accurate pose annotations, which are usually obtained by running SfM on densely captured images [21] or human labeling [7].
+As it is expensive to collect such dense data at large scale, the generalization ability of these methods is often limited.
+In fact, current experiments are primarily conducted on turntable-like videos of common objects [21], which exhibit a relatively simplistic appearance and viewpoint distributions in comparison to that of open-world images.
+We solve the sparse-view camera pose estimation problem by harnessing the pre-trained diffusion model of novel views conditioned on viewpoints (Zero-1-to-3) [13].
 We present ID-Pose which inverses the denoising diffusion process to estimate the relative pose of two input images.
-ID-Pose leverages the pre-trained diffusion model, and maintains the relative pose as the optimization variable.
-It is conditioned on one image and the pose to predict the noise added to the other image.
+ID-Pose leverages the noise predictor of the diffusion model, and maintains the relative pose as the decision variable.
+The predictor conditions on one image and the current value of the pose to predict the noise that is added to the latent of the other image.
 The error of the prediction is used as the objective to update the pose with the gradient descent method.
-To prevent local minimum, we propose an initialization strategy that first samples a number of candidature poses with a few steps of optimization and then selects the one with the minimum prediction error of noise.
+To improve the stability, we propose a two-stage strategy including an exploration process that finds a promising initial pose, and then a refinement process that further enhances the accuracy of the pose.
+The exploration process first samples a number of candidature poses with a few rounds of optimization and then selects the one with the minimum prediction error of noise.
+The refinement process keeps updating the selected pose until the stopping criteria are satisfied.
 We further extend ID-Pose to effectively find relative poses of more than two images.
@@
 Within a group of three images, we can use the poses of two image pairs to present the pose of the third pair.
-Therefore, the prediction error of the third pair can be used to update the two poses alternately.
-ID-Pose is a zero-shot method that generalizes to open-world images by leveraging the diffusion model pre-trained on large-scale data [8, 16].
-We conduct experiments on high-quality real-scanned 3D objects, where ID-Pose significantly outperforms state-of-the-art methods.
+Therefore, the prediction error of the third pair can be used to update the two poses.
+This triangular relation enables the incorporation of additional information, thereby improving the accuracy of the estimation.
+In general, ID-Pose transfers the image generation ability from the large-scale model to the camera pose estimation ability.
+The advantages of the method are as follows:
+  • It is a zero-shot method that requires no training;
+  • Generalizes to open-world images with varying object categories, image styles, and viewpoints;
+  • Based on the understanding of objects, so that appearance overlaps are not required in input views.
+We conduct extensive experiments on NAVI [7], OmniObject3D [30], ABO [2], and CO3D [21] datasets, which include both casually captured photos and rendered images with randomly sampled viewpoints.
+The results demonstrate that ID-Pose outperforms state-of-the-art methods with strong generalization capability.
 2 Related Work
 Camera Pose Estimation.
-Estimating camera poses of images is a long-standing problem.
-Traditional methods extract local features [9, 19, 1] of input images and find their matches to estimate poses.
-When images are sparsely captured, they often fail to estimate poses due to insufficient feature matches.
-Recent learning based methods [10, 22, 18, 7] use neural networks to predict the relative poses of image pairs.
-These methods require a large number of pose annotations, which are usually estimated with dense views of an object [15].
-As training data is expensive to obtain, the neural networks are limited to benchmarks and hard to handle out-of-distribution images.
+Estimating the camera poses of images is a fundamental problem in 3D computer vision.
+Traditional methods [26] extract local features [14, 28, 1] of input images and find feature correspondences to estimate the camera poses.
+When images are sparsely captured, they often fail due to insufficient feature matches.
+Recent learning-based methods [15, 35, 27, 10] use neural networks to predict the relative poses of image pairs.
+These methods allow sparse-view input, but the training requires a large number of pose annotations, which are usually obtained from dense views [21].
+As training data is expensive to acquire, the neural networks are limited to benchmarks and hard to handle out-of-distribution data.
 Image Diffusion Models.
 Diffusion models [6] have been widely used for image generation tasks.
-They learn a neural network to denoise an image, that takes a noisy image to predict the noise values mixed in the image.
-As the noise is random, the model essentially predicts the image and subtracts it from the noisy image to obtain the noise values.
-By progressive denoising from pure noise, a clear image can be generated.
-Latent diffusion models [16] denoise latent representations of images, which reduce the model complexity and improve the efficiency and stability.
-Recent models are able to condition on other inputs to control the generation, for example, natural language and sketches [23, 12].
+These models learn a neural network to denoise a noisy image, that predicts the noise value added to the image and subsequently subtracts it.
+By progressive denoising from pure noise, a clear and complex image can be generated.
+Latent diffusion models [22] denoise latent representations of images, which improves efficiency and stability.
+Recent models are able to control the generation conditioned on other inputs, for example, natural language and sketches [36, 18].
 Stable Diffusion [22] is a latent diffusion model that generates high-quality images based on input texts by training on over a hundred million image-text pairs.
-A view-conditioned diffusion model, Zero-1-to-3 [8], uses one input image of an object and a relative pose to generate the novel view of the object under the new viewpoint.
-It is fine-tuned from Stable Diffusion with pose-annotated images rendered using Objaverse dataset [3].
-It achieves prominent generation results with real-world images.
-We consider to inverse the denoising process of Zero-1-to-3, that uses two images of the same object to predict their relative camera pose.
-This enables zero-shot pose estimation given a sparse set of images.
+A view-conditioned diffusion model, Zero-1-to-3 [13], uses an input image of an object and a relative pose to generate a novel view of the object under the new viewpoint.
+It is fine-tuned from Stable Diffusion with pose-annotated images rendered from the Objaverse dataset [3], and achieves prominent performance on real-world images.
+Given that Zero-1-to-3 associates views with viewpoints, we explore the possibility of utilizing this model to reveal the relative camera pose between two images.
+This will enable zero-shot camera pose estimation in the context of sparse views with small or no overlapping appearance.
+Inversion Techniques Given a feed-forward model that maps an input to the corresponding output, inversion techniques convert the output back to the input using optimization.
+In image editing, inversion is commonly used to convert an image to a latent vector, and one can edit the vector to modify the image [38, 31].
+By using Stable Diffusion, TextualInversion [5] inverts images to text embedding that enables the generation of new images with the same concepts or objects.
+iNeRF [33] showcases that the camera pose of an image can be inversely found conditioned on a neural 3D scene [16].
+Inspired by these approaches, we investigate inversely using novel-view generative models to estimate relative camera poses of given images, through optimization-based techniques.
 3 Methodology
@@
 We present ID-Pose, a zero-shot method that takes two images as input and finds their relative camera pose.
-The idea is to invert the denoising process of the pre-trained novel view generation model, i.e., Zero-1-to-3 [8].
-We extend ID-Pose to handle more than two images by inverting each of the poses with multiple image pairs.
+The idea is to invert the denoising process of the pre-trained novel-view diffusion model, i.e., Zero-1-to-3 [13], where the error of the noise prediction is minimized to obtain the pose.
+We further extend ID-Pose to handle more than two images by inverting each of the poses with multiple image pairs.
 In this section, we first introduce the background knowledge of diffusion models, and then describe the details of the ID-Pose method.
@@
 Zero-1-to-3 [13] is fine-tuned from Stable Diffusion that generates novel views of an object conditioned on a reference view $x_{c}$ and a relative camera pose $p$.
-Its noise predictor is written as $f(z(\epsilon,t),x_{c},p,t)$.
+The noise predictor of Zero-1-to-3 is written as $f(z(\epsilon,t),x_{c},p,t)$.
+For convenient representations, we simplify this notation to $f(z(\epsilon),x,p)$ where $t$ is omitted.
 The reference view $x_{c}$ is mapped to image tokens with the CLIP image encoder.
-The camera pose $p$ is expressed in a spherical coordinate system where the position of a point is specified by three numbers: polar angle, azimuth angle, and radius.
+The camera pose $p$ is expressed in the spherical coordinate system where the position of a point is specified by three numbers: polar angle, azimuth angle, and radius.
 The three values of the pose $p$ are concatenated with the $x_{c}$ tokens as conditions to the predictor $f$.
-Zero-1-to-3 is fine-tuned with large-scale images rendered from 3D objects but generalizes to real-world images.
+Zero-1-to-3 is fine-tuned with large-scale images rendered from 3D objects [3] but generalizes to real-world images.
 3.2 ID-Pose
-Let $f(z(\epsilon),x,p)$ be the pre-trained noise predictor of Zero-1-to-3, where the diffusion step $t$ is omitted for simple notations.
-Given two images $x_{0}$ and $x_{1}$, we inverse the denoising process to estimate their relative pose, which is represented in spherical coordinates (polar, azimuth, radius).
-Let $x_{0}$ be the reference view and $x_{1}$ be the target view.
-We add a noise $\epsilon$ to $x_{1}$ and use $f$ to predict the noise.
-Assume $p$ is the pose transformation from $x_{0}$ to $x_{1}$, the predicted noise is defined as a function of $p$:
+Recall that $f(z(\epsilon),x,p)$ is the pre-trained noise predictor of Zero-1-to-3.
+Given two images $x_{0}$ and $x_{1}$, we want to use $f$ to find their relative pose $p$ that is represented in spherical coordinates (polar, azimuth, radius).
+Assume $x_{0}$ is the reference view and $x_{1}$ is the target view.
+We can add a noise $\epsilon$ to $x_{1}$ and use $f$ to predict that noise.
+We define $p$ as the pose transformation from $x_{0}$ to $x_{1}$, the predicted noise can be considered as a function of $p$:
 $\hat{\epsilon}(p\,;\,x_{0},x_{1})=f(z_{1}(\epsilon),x_{0},p),$ (1)
@@
 We also consider that $x_{1}$ acts as the reference view and $x_{0}$ acts as the target view.
-Let $p^{\prime}$ be the reversed transformation of $p$ (from $x_{1}$ to $x_{0}$).
-The predicted noise of $p^{\prime}$ is:
+Let $p^{\prime}$ be the reversed transformation of $p$ (from $x_{1}$ to $x_{0}$), where $p^{\prime}=-p$ is held in spherical coordinates.
+The predicted noise regarding $p^{\prime}$ is defined as:
 $\hat{\epsilon}(p^{\prime}\,;\,x_{1},x_{0})=f(z_{0}(\epsilon),x_{1},p^{\prime}),$ (3)
-where $z_{0}(\epsilon)$ is the noisy latent map of $x_{0}$. The noise error of the prediction given $p^{\prime}$ is:
+where $z_{0}(\epsilon)$ is the noisy latent map of $x_{0}$. Thereby, the noise error of the prediction given $p^{\prime}$ is:
 $l(p^{\prime}\,;\,x_{1},x_{0})=||\hat{\epsilon}(p^{\prime}\,;\,x_{1},x_{0})-\epsilon||_{2}^{2}.$ (4)
-We randomly choose $l(p\,;\,x_{0},x_{1})$ or $l(p^{\prime}\,;\,x_{1},x_{0})$ as the objective function $g(p)$. The inversion process iteratively updates $p$ by the gradient descent method:
-$p^{(k+1)}=p^{(k)}-\alpha\nabla g(p^{(k)})$ (5)
-where $\alpha$ is the step factor, $k$ represents the current updating step, and $k+1$ represents the next step.
-The iteration stops when $p$ converges or $k$ exceeds a maximum step count.
-The intuition is that, $p$ reaches the optimum once the model can predict the noise with the minimum error.
-The initialization of $p$ is crucial as optimization can be often stuck into local minimums.
-For a near-symmetric object, the opposite view of one view can have the highest similarity among the other views.
-Under this situation, if $p$ is initialized near the opposite view, it will be likely stuck into a local minimum.
-To alleviate this problem, we use a sampling approach to find reliable initialization.
-We uniformly sample a number of pose candidates on a sphere around the origin point.
-The relative radius of the candidature poses is set to zero.
-For each candidate, we update the pose for a few steps with Equation 5.
-We then evaluate those poses with the pairwise error:
+We can inversely find $p$ by minimizing either of the noise errors:
+$\min_{p}{l(p\,;\,x_{0},x_{1})\;\text{or}\;l(p^{\prime}\,;\,x_{1},x_{0})}.$ (5)
+We alternately choose $l(p\,;\,x_{0},x_{1})$ or $l(p^{\prime}\,;\,x_{1},x_{0})$ as the objective function $g(p)$. The optimization process iteratively updates $p$ using the gradient descent method:
+$p^{(k+1)}=p^{(k)}-\alpha\nabla g(p^{(k)}),$ (6)
+where $\alpha$ is the updating factor, $k$ represents the current iteration, and $k+1$ represents the next iteration.
+The optimization stops when $p$ converges or $k$ exceeds a preset maximum count.
+The intuition is that, $p$ reaches the optimum once the model $f$ predicts the noise with the minimum error.
+The initial value of $p$ is crucial as optimization can be often stuck into local minimums.
+In the case of nearly symmetric objects, the opposing view of a targeting view may exhibit the greatest similarity when compared to the other views.
+Under this situation, if $p$ is initialized near this opposing view, it will be likely stuck into a local minimum.
+To alleviate this problem, we design ID-Pose with two stages.
+The first stage runs an exploration process to find a reliable initial value of $p$.
+The second stage runs a refinement process to further update $p$ with Equation (6) until stops.
+We describe the exploration process as follows.
+We first create $m$ candidature relative poses, where the azimuth angles are uniformly sampled from 00 to $2\pi$.
+We then use the method from [12] to estimate the absolute elevations of the two images, where the difference is used to set the relative polar angles of the candidates.
+We lastly set the relative radius to zero.
+For each candidature pose, we update it for a few iterations with Equation (6).
+We then probe those poses with the pairwise error:
 $l_{0,1}(p)=l(p\,;\,x_{0},x_{1})+l(p^{\prime}\,;\,x_{1},x_{0}),$ (7)
-which combines the two noise errors to measure the stability of $p$ connecting $x_{0}$ and $x_{1}$.
-From all the candidates, we pick the one with the minimum pairwise error as the initial pose and further update it until convergence.
+which combines the two noise errors to measure the instability of $p$ connecting $x_{0}$ and $x_{1}$. From all the candidates, we pick the one with the minimum pairwise error as the initial pose.
 3.3 ID-Pose with More Views
-Given $n$ ($n>2$) images, a naive strategy is breaking them into $n-1$ consecutive pairs, and inverting the pose for each of them independently.
-However, there are many unused image pairs that can help to find these poses.
-We consequently propose to estimate each of the relative poses with multiple image pairs from triangular relationships.
+Given more than two images, i.e., $n>2$, a naive strategy is selecting an anchor image and creating $n-1$ image pairs, where each pose is inverted independently.
+However, there are many connectable but unused relations that can help find these poses jointly.
+We consequently consider estimating each of the relative poses with multiple image pairs from triangular relations.
 Suppose the input images are $x_{0}$, $x_{1}$, …, $x_{n-1}$, we maintain $n-1$ relative poses $p_{1}$, $p_{2}$, …, $p_{n-1}$ as variables, where $p_{i}$ represents the relative transformation from $x_{0}$ to $x_{i}$.
-Firstly, we describe the initialization process.
-For each $p_{i}$, we uniformly sample $m$ candidates and update them with a few steps.
+Firstly, we describe the exploration process.
+For each $p_{i}$, we uniformly sample $m$ candidates and update them with a few iterations.
 Instead of picking the best one, we keep those candidates as a group $\{p_{i,u}|u\in[0,m)\}$.
@@
 where $p^{\prime}_{i}+p_{j}$ is the relative transformation from $x_{i}$ to $x_{j}$.
-The triangular error sums up three pairwise errors, which indicates the instability of poses $p_{i}$ and $p_{j}$ to connect views $x_{0}$, $x_{i}$, $x_{j}$.
+The triangular error sums up three pairwise errors, which indicates the instability of poses $p_{i}$ and $p_{j}$ to connect the three views $x_{0}$, $x_{i}$, and $x_{j}$.
 Then, given a candidate $p_{i,u}$ of $p_{i}$, we iterate the candidates $p_{j,v}$ of $p_{j}$ and calculate triangular errors ($m$ times).
@@
 $e_{i,u,j}=\min_{v\in[0,m)}l_{i,0,j}(p_{i,u},p_{j,v}).$ (9)
-We obtain $e_{i,u}=\sum_{j\neq i}e_{i,u,j}$ as the instability of $p_{i,u}$. We select $p_{i,u}$ with lowest instability as the initialization of $p_{i}$.
-Next, we describe the inversion process.
+We obtain $e_{i,u}=\sum_{j\neq i}e_{i,u,j}$ as the instability of $p_{i,u}$.
+We select $p_{i,u}$ with lowest instability as the initialization of $p_{i}$.
+Next, we describe the refinement process.
 Suppose all poses are initialized, we randomly pick two indices $i$ and $j$ where $0\leq i\neq j<n$.
@@
 We fix one and update the other with the noise error $l(p^{\prime}_{i}+p_{j}\,;\,x_{i},x_{j})$.
+The refinement process stops upon reaching a preset maximum iteration count.
 4 Experiments
 4.1 Experimental Settings
-Data.
-We conduct experiments on two datasets: OmniObject3D [20] and Amazon Berkeley Objects (ABO) [2].
-These datasets are only used for testing, and none of the experimental methods are trained on them.
-OmniObject3D is a dataset of high-quality real-scanned 3D objects.
-It includes 190 categories with 6,000 scanned objects.
+Data Preparation.
+We conduct experiments on 4 datasets: NAVI [7], OmniObject3D [30], Amazon Berkeley Objects (ABO) [2], and Common Objects in 3D V2 (CO3D) [21].
+As the optimization in ID-Pose requires a certain amount of time, we pick a number of samples from each dataset to create testing sets.
+NAVI [7] is a dataset of casually captured images of 36 category-agnostic objects using hand-held cameras.
+For each object, we randomly select one set of multi-view images.
+Each set of images is used as a view pool, where we randomly pick out images for pose estimation evaluations.
+OmniObject3D [30] is a dataset of high-quality real-scanned 3D objects with 190 categories.
 We select 12 categories “house”, “chair”, “table”, “bed”, “ornaments”, “shoe”, “scissor”, “guitar”, “laundry detergent”, “dumbbell”, “dinosaur”, and “glasses”.
-We select the first 5 scanned objects for each category if available.
-The total number of objects is 54.
-For each object, we use all the Blender rendered images as a view pool, and change their background color to white.
-ABO is a dataset of 3D Amazon.com products.
-We select 10 representative objects: “bag”, “bar stool chair”, “ergonomic chair”, “sofa”, “couch”, “desk”, “shelf”, “bed”, “fan”, and “lamp”.
-For each object, we place its geometry center at the origin point.
-We then setup a camera with FOV=$60^{\circ}$ looking at the origin, and randomly place it on a sphere to render 30 images.
-The radius of the sphere is set to a value that the object can be fully captured with enough image margin.
-The background is set to white during rendering.
-The rendered images form a view pool of an object.
-The ground truth poses are known and associated with the images.
-Given a view pool, we randomly pick out $n$ images as a testing sample.
-We experiment with $n$ = $2,4,6,8$.
-For OmniObject3D, we repeat the picking 3 times.
-We make $648$ testing samples in total containing 2,592 relative poses to estimate.
-For ABO, we repeat the picking 10 times.
-We make $400$ testing samples in total containing 1,600 relative poses to estimate.
+For each category, we select the first 5 scanned objects if available.
+We totally select 54 objects.
+For each object, we use their Blender rendered images (modified to white background) as a view pool.
+ABO [2] is a dataset of 3D synthetic objects of commercial products on Amazon.com.
+We select 10 objects: “bag”, “bar stool chair”, “ergonomic chair”, “sofa”, “couch”, “desk”, “shelf”, “bed”, “fan”, and “lamp”.
+For each object, we render 30 images to create a view pool.
+We place the object at the origin point.
+We then setup a camera with FOV=$60^{\circ}$ looking at the origin.
+We randomly move cameras on a sphere to render images.
+The background is set to white during the rendering.
+CO3D (V2) [21] is a dataset of turntable-style video sequences captured around real-world objects.
+It includes 51 object categories.
+We select 12 categories “bench”, “bicycle”, “car”, “chair”, “hydrant”, “parkingmeter”, “teddybear”, “toaster”, “toybus”, “toyplane”, “toytruck”, “toytrain”.
+For each category, we select 5 sequences from the test split.
+For each sequence, we extract every 10 consecutive frames, which are used as a view pool.
+Given a view pool, we pick out images to make testing samples.
+We define the pick operation as randomly choosing 8 images from the pool.
+We order the images, and select the first 2, 4, 6, 8 images to make four testing samples, respectively.
+This eliminates data variants in comparison between using less and more views.
+For NAVI, we make 210$\times$4 testing samples including 3360 pairs of images to estimate poses.
+For OmniObject3D, we make 162$\times$4 samples including 2592 pairs.
+For ABO, we make 100$\times$4 samples including 1600 pairs.
+For CO3D, we make 180$\times$4 samples including 2880 pairs.
+Given a sample with $n$ images, we use the first one as the anchor view and $n-1$ target views.
 Baseline Settings.
-We use RelPose++ [7] as the baseline method, which is trained using Common Objects 3D (v2) dataset [15].
+We compare ID-Pose with state-of-the-art methods: 1) COLMAP: The Structure-from-Motion pipeline that uses SuperPoint [4] features with LightGlue [11] matching.
+We use the implementation provided by HLoc [24, 25].
+2) RelPose++ [10]: RelPose++ is the state-of-the-art method that uses network to predict 6DOF camera poses.
 There are two released model checkpoints.
-One is trained using raw images, the other is trained using masked images with clean background.
-We evaluate both of them in the experiments.
+One is trained using full images, the other is trained using masked images where background is removed.
+We evaluate both of them in the experiments, and separately denote them as FULL and MASK.
 ID-Pose Settings.
 We use the checkpoint of Zero-1-to-3 [13] trained with 105000 iterations.
-We preprocess input images to have white background and contain enough margin, which fit the input distribution of the noise predictor.
-Given an image, we use a segmentation mask to clear the background, and use a window to crop out the object at its center.
+We preprocess input images to feature a white background and sufficient margins, thereby aligning with the input distribution of the noise predictor.
+Given an image, we use a segmentation mask to clear the background and use a window to crop out the object at its center.
 The size of the window is initially set to the largest bounding box of segmentation masks of all input images.
 The size is then scaled 1.5x to ensure that there is enough margin in the cropped images.
-The hyper-parameters used in ID-Pose are shown in Table 1.
-We additionally test ID-Pose without using triangular relationships.
-Initialization Stage Optimization Stage
-Noise $\epsilon$ fixed random (each iteration)
-Noise level 0.8 0.8
-Step factor $\alpha$ 10.0 1.0
-Batch size 16 1
-Iterations 10 (each pose) $600\times(n-1)$
-Table 1: The hyper-paramters of ID-Pose.
-Metrics.
-Given $n$ images and the camera pose of one image.
-We estimate relative poses and obtain the poses for the other $n-1$ images.
-The poses are transformed into world coordinate system.
-Following RelPose++[7], we obtain relative rotations of the estimated poses and compute the angular differences to the ground truth relative rotations.
+The parameters used in ID-Pose are shown in Table 1.
+They are used across all datasets.
+We additionally evaluate two ablated variants: 1) ID-Pose Naive: finding poses independently for each pair of the images; 2) ID-Pose TR: exploring poses independently and using triangular relations only during the refinement process.
+Exploration Refinement
+Probe (7) Update Update
+Noise $\epsilon$ fixed random random
+Noise step 0.2 [0.2, 0.8] [0.2, 0.8]
+Factor $\alpha$ 10.0 1.0
+Batch size 16 1 1
+Iterations - 10 (each pose) $600\times(n-1)$
+Table 1: The parameters used in ID-Pose.
+Evaluation Settings & Metrics.
+Given $n$ images and the camera pose of a selected anchor view.
+We estimate the relative poses and obtain the absolute poses for the other $n-1$ images.
+To obtain the absolute camera positions, we scale the relative position with the radius of the anchor view position.
+The poses are compared with the ground truth in the world coordinate system.
+Following RelPose++[10], we obtain relative rotations of the estimated poses and compute the angular differences to the ground truth.
 We report the rotation accuracy which is the proportion of the angular differences less than 15 or 30 degree.
@@
 $n=2$ $n=4$ $n=6$ $n=8$
-Method $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA
-RelPose++ w/o Mask 5.56 14.20 8.64 4.12 12.35 7.20 3.70 12.72 8.27 4.41 12.61 9.08
-RelPose++ 6.17 11.11 12.96 4.32 11.73 12.76 3.70 11.85 11.60 4.50 11.90 11.46
-ID-Pose w/o Tri 48.77 56.79 45.06 48.77 59.88 44.65 48.77 60.25 44.57 49.65 61.38 45.06
-ID-Pose 48.77 56.79 45.06 49.79 61.11 46.50 50.12 63.33 46.42 52.38 65.26 48.06
-Table 2: Results on the OmniObject3D testing samples.
-“$15^{\circ}$”: rotation accuracy by $15^{\circ}$.
-“$30^{\circ}$”: rotation accuracy by $30^{\circ}$.
+$15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA
+NAVI Pairs 210 Pairs 630 Pairs 1050 Pairs 1470
+COLMAP 25.24 26.67 0.00 16.35 16.51 0.16 12.67 12.76 0.10 11.09 11.16 0.41
+$\vdash$ Converge # 81 + 0 of 210 32 + 62 of 210 11 + 85 of 210 11 + 87 of 210
+RelPose++ FULL 38.10 63.33 21.43 36.03 62.38 21.90 35.71 61.62 22.95 36.67 62.93 22.52
+RelPose++ MASK 25.24 51.43 17.14 25.08 49.52 15.87 24.10 49.71 15.90 22.45 48.37 13.88
+ID-Pose Naive 48.57 81.43 42.38 53.33 79.52 44.29 52.10 79.24 44.10 53.20 79.18 43.88
+ID-Pose TR 48.57 81.43 42.38 56.35 83.49 47.78 58.29 83.24 48.19 60.61 83.20 50.27
+ID-Pose 48.57 81.43 42.38 57.30 85.56 48.25 60.19 86.76 50.48 62.52 87.62 51.84
+OmniObject3D Pairs 162 Pairs 486 Pairs 810 Pairs 1134
+COLMAP 12.35 15.43 0.00 8.23 9.88 0.00 8.27 9.63 0.37 8.73 9.88 0.88
+$\vdash$ Converge # 53 + 0 of 162 20 + 43 of 162 16 + 51 of 162 19 + 51 of 162
+RelPose++ FULL 12.35 30.25 8.64 9.47 27.16 7.20 9.88 27.16 8.27 10.93 30.51 9.08
+RelPose++ MASK 16.05 38.27 12.96 13.37 34.36 12.76 12.72 33.83 11.60 12.17 33.33 11.46
+ID-Pose Naive 70.99 86.42 62.35 64.81 80.66 54.73 64.44 80.25 55.06 63.49 79.81 55.11
+ID-Pose TR 70.99 86.42 62.35 69.34 82.10 61.52 70.49 82.84 62.59 69.05 83.16 63.76
+ID-Pose 70.99 86.42 62.35 67.70 81.07 59.47 71.23 84.07 64.20 70.11 84.13 64.02
+ABO Pairs 100 Pairs 300 Pairs 500 Pairs 700
+COLMAP 5.00 7.00 0.00 2.00 2.67 0.00 1.20 1.60 0.00 1.71 2.00 0.00
+$\vdash$ Converge # 12 + 0 of 100 3 + 10 of 100 0 + 13 of 100 2 + 12 of 100
+RelPose++ FULL 4.00 9.00 2.00 3.33 8.00 1.67 3.80 9.80 3.20 3.57 10.57 3.43
+RelPose++ MASK 3.00 8.00 1.00 2.33 9.67 1.67 4.20 9.60 4.00 4.29 10.00 3.86
+ID-Pose Naive 52.00 64.00 43.00 47.67 64.33 41.33 46.60 60.80 40.80 50.71 64.86 43.86
+ID-Pose TR 52.00 64.00 43.00 53.00 64.67 45.33 52.80 62.00 46.80 57.86 65.71 50.29
+ID-Pose 52.00 64.00 43.00 52.00 61.00 44.33 55.00 63.40 49.00 59.29 68.00 53.00
+CO3D Pairs 180 Pairs 540 Pairs 900 Pairs 1260
+COLMAP 20.00 25.56 0.00 19.44 21.67 0.93 22.44 24.11 2.00 23.89 26.51 3.25
+$\vdash$ Converge # 86 + 0 of 180 48 + 54 of 180 54 + 57 of 180 60 + 60 of 180
+RelPose++ FULL 83.33 92.22 47.22 84.63 94.44 49.44 85.78 94.11 51.11 86.83 94.52 50.24
+RelPose++ MASK 81.67 90.00 46.67 82.41 91.30 49.63 81.44 91.22 49.33 82.22 91.59 47.38
+ID-Pose Naive 39.44 64.44 32.22 42.22 66.30 32.59 42.00 69.67 32.78 41.51 67.30 31.83
+ID-Pose TR 39.44 64.44 32.22 46.30 69.07 36.48 46.44 72.78 40.56 46.59 71.11 38.65
+ID-Pose 39.44 64.44 32.22 47.22 70.56 37.96 46.56 73.56 39.89 48.49 72.70 39.29
+Table 2: Results on the testing samples of all datasets.
+“$15^{\circ}$”: rotation accuracy with $15^{\circ}$ threshold.
+“$30^{\circ}$”: rotation accuracy with $30^{\circ}$ threshold.
 “PA”: camera position accuracy.
-The numbers are shown in 100%.
-$n=2$ $n=4$ $n=6$ $n=8$
-Method $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA
-RelPose++ w/o Mask 2.00 6.00 2.00 2.00 7.33 1.67 3.00 7.40 3.20 2.86 7.86 3.43
-RelPose++ 2.00 6.00 1.00 2.00 7.00 1.67 4.00 7.60 4.00 3.43 7.29 3.86
-ID-Pose w/o Tri 76.00 80.00 73.00 75.00 78.33 74.00 77.60 80.00 75.60 77.71 80.57 75.86
-ID-Pose 76.00 80.00 73.00 77.00 79.33 74.00 78.00 81.40 75.60 78.71 82.14 77.14
-Table 3: Results on the ABO testing samples.
-We report evaluation results aggregated by the number of input views.
-The results on OmniObject3D and ABO samples are shown in Table 2 and Table 3, respectively.
-ID-Pose methods significantly outperform RelPose++ for all different $n$, which demonstrate advanced generalization capability.
-RelPose++ methods report much lower metrics compared to the testing results on the CO3D dataset [7].
-The reason might be that the models of RelPose++ overfit to CO3D data and fail to generalize to out-of-distribution images.
-With two input images ($n=2$), ID-Pose w/o Tri and ID-Pose report the same results as they behave identically.
-With more input images ($n>2$), ID-Pose outperforms ID-Pose w/o Tri as extra image pairs are used to estimate poses.
-By adding input images, the performance of ID-Pose increase as more image pairs are used.
-We visualize examples of estimated camera poses in the world coordinate system.
-We compare ID-Pose and Relpose++ with ground truth.
-Assume that the true camera pose of the reference image is known.
-For RelPose++, we transform all predicted poses with a matrix that aligns the first prediction to the reference.
-In Figure 2, we input four views of the “house_001” from OmniObject3D, where the front view acts as the reference.
-Compared with the ground truth, ID-Pose finds all three relative poses with small errors.
-RelPose++ accurately predicts the relative poses of the two side views, but makes a large prediction error on the opposite view.
-In Figure 3, we input four views of the “ornaments_001” from OmniObject3D, where the “left-face” view acts as the reference.
-This is a hard case for both ID-Pose and RelPose++.
-ID-Pose finds the approximate poses but with large errors compared to the ground truth.
-RelPose++ predicts poses with even larger errors.
-In Figure 4, we input four views of an ABO chair object.
-ID-Pose accurately finds the camera poses, while RelPose++ fails to distinguish the differences between these images.
-This failure happens to almost all the ABO testing samples.
-RelPose++ is probably hard to handle out-of-distribution images.
-This exposes the weakness of the fitting based methods, that the generalization capability is limited by the amount of annotated training data.
-ID-Pose leverages the model pre-trained on large-scale images, which demonstrates better generalization ability.
-Figure 2: Visualization of camera poses with a sample of views from OmniObject3D “house_001”.
-Figure 3: Visualization of camera poses with a sample of views from OmniObject3D “ornaments_001”.
-Figure 4: Visualization of camera poses with a sample of views from an ABO chair.
-4.3 Qualitative Results on In-the-wild Images
-We qualitatively show the results of ID-Pose on in-the-wild images.
-We find three objects in our daily workplace: a toy duck, a chair, and a foosball table.
-We use a smartphone to capture 4-6 images for each of them.
-We use the DIS method [13] to segment out the objects and clean the background of the images to white.
+The values of the results are shown in 100%.
+[…]
+We report evaluation results aggregated by the number of input views $n$.
+The results on four datasets are shown in Table 2.
+We compare ID-Pose with the state-of-the-art RelPose++.
+According to the results on NAVI, OmniObject3D, and ABO samples, ID-Pose significantly outperforms RelPose++ for all different $n$.
+ID-Pose demonstrates advanced generalization capability on both casually captured photos and rendered images from arbitrary viewpoints.
+For CO3D samples, ID-Pose shows lower accuracy than RelPose++, which exhibits remarkable performances on this dataset.
+The reason might be that RelPose++ is trained with CO3D data and the testing samples share a similar distribution.
+However, as RelPose++ is based on fitting neural networks, the method could fail with out-of-distribution images and camera poses.
+In comparison, ID-Pose does not require training and generalizes well to a wide range of images.
+Note that RelPose++ FULL performs better than MASK on CO3D and NAVI, while performs worse on OmniObject3D and ABO.
+The reason is that the images of CO3D and NAVI include environmental background.
+This provides extra information which is learned by FULL to improve accuracy.
+The images of OmniObject3D and ABO has no background, which are out-of-distribution data for FULL.
+We find that COLMAP frequently fails to converge with sparse views, due to lack of reliable features from views with large appearance differences.
+We additionally report converge counts of running COLMAP.
+The counts includes 3 numbers A + B of C.
+A is the count of successful convergences with all $n$ images.
+B is the count of sub-success cases that the running with all images failed but there exists a success with fewer images.
+C is the total number of the samples.
+The successful count is higher for $n=2$, as joining two views is simple.
+With $n>2$, solving joint matches of multiple sparse views becomes difficult.
+We compare ID-Pose with the two ablated variants: Naive and TR.
+With $n=2$, they show the same results as there is no triangular relation.
+With $n>2$ , ID-Pose mostly outperforms the others, and TR outperforms the Naive.
+This demonstrates the effectiveness of using triangulation.
+There are some exceptions with $n=4$ for OmniObject3D and ABO, where TR performs the best.
+As these images are rendered from random viewpoints, there could be abnormal views that affect the exploration process to find promising initial poses.
+We visualize examples of estimated camera poses in the world coordinate system, under the assumption that camera poses of anchor views are known.
+We compare ID-Pose with RelPose++ (best version).
+In Figure 2, we show two examples from NAVI and two examples from OmniObject3D, where the cameras are presented by 3D cones.
+The rotations and shifts between the estimated cones and the ground truth (grey color) can be observed.
+ID-Pose (blue color) accurately finds the camera poses, and surpasses the performance of RelPose++ (green color).
+As these objects are category-agnostic, ID-Pose shows better generalization ability due to leveraging the generative model pre-trained on large-scale images.
+Note that the bottom-right sample is a “dumbbell” which exhibits symmetry along its central axis.
+It is intractable to differentiate two symmetrically opposite views, e.g., image 1 and image 2, based on pure object appearance.
+However, ID-Pose conditions on lighting to find accurate camera poses.
+In Figure 3, we show two examples from ABO and two examples from CO3D.
+Given ABO samples, which are synthetic objects, ID-Pose yields accurate results as expected.
+But RelPose++ fails to predict the poses as it cannot generalize to synthetic content.
+Given CO3D samples, ID-Pose under-performs RelPose++ which might overfit to this particular data distribution.
+In general, ID-Pose based on iterative optimization finds more accurate poses than feed-forward prediction.
+We show some failed cases of ID-Pose.
+In the “chair” example with $n=4$, the pose for image 3 is incorrect.
+In the “bicycle” example with $n=6$, the poses of the image 2 and 5 are incorrect.
+We find that these prediction are axis-symmetric to the ground truth.
+This indicates that similar symmetric views can impact the performance of ID-Pose.
+[…]
+4.3 Noise Step Study
+[…]
+4.4 Using Self-captured Images
+We qualitatively show the results of ID-Pose on self-captured images.
+We find four objects in our daily workplace: a toy duck, a toy car, a chair, and a foosball table.
+We use a smartphone to capture 4-6 images for each of them, and use the DIS method [19] to segment out the objects with clean background.
 We select $n$ ($n=2,3,4$) images as input views to estimate the relative poses.
 To visualize the results, we use the LumaAI APP [^1] https://lumalabs.ai/ to reconstruct 3D meshes of the objects.
-We then choose one image as the reference and manually find its camera pose using the 3D mesh.
-With the ground truth camera pose of the reference image, we transform the estimated poses of the target views into the world coordinate system.
+We then choose one image as the anchor and manually find its camera pose using the 3D mesh.
+With the ground truth camera pose of the anchor image, we transform the estimated poses of the target views into the world coordinate system.
 We show an example of the results of the chair in Figure 5.
-More results can be interactively visualized on the project website [^2] https://xt4d.github.io/id-pose/.
-The estimated poses cannot align the views to the 3D object perfectly.
-One reason is that the true cameras may not look at the same point in the world, and their “up” vectors can be varied.
-For in-the-wild images, pose estimation with the assumption of spherical coordinates will introduce additional errors.
-This is a weakness of ID-Pose.
-Figure 5: Visualization of estimated camera poses of sparsely captured views around a chair.
+More results are provided in our Project Page.
+We observe that the estimated poses cannot align the views to the 3D object perfectly.
+A reason is that the cameras may not look at the same point in the world, and their “up” vectors can be varied.
+Therefore, pose estimation with the assumption of spherical coordinates will introduce additional errors.
+4.5 GPU Memory Usage & Running Time
+We conduct experiments using a single NVIDIA V100 GPU 32G.
+Running the method requires loading a Zero-1-to-3 checkpoint, where the GPU memory usage is about 28G.
+The running time includes exploration time and refining time.
+For two images, the exploration time is about 1 minute.
+The refining takes about 30 seconds per 100 iterations.
+Running 600 iterations takes about 3 minutes.
+Figure 4: The metrics with respect to noise step $t$.
+Figure 5: Visualization of the estimated camera poses of the captured images around a chair.
 The mesh is reconstructed using LumaAI APP.
-The view in red acts as the reference view, where the camera pose is manually found using the mesh.
-5 Conculsion & Limitations
+The view in red acts as the anchor view, where the camera pose is manually found using the mesh.
+5 Conclusion & Limitations
 In this paper, we propose ID-Pose which inverses the denoising diffusion process to estimate camera poses of sparse input views.
-ID-Pose only utilizes the pre-trained Zero-1-to-3 model [8] without additional training.
-Given two images, ID-Pose adds a noise on one image, and predicts the noise conditioned on the other image and a decision variable for the relative pose.
-The prediction error is used as the objective to find the optimal pose with the gradient descent method.
+ID-Pose is a zero-shot method based on the pre-trained Zero-1-to-3 model [13].
+To find the relative pose of two given images, ID-Pose uses the noise predictor, which conditions on one image and a hypothesis of the relative pose, to predict the noise added on another image.
+The prediction error is used as the objective function to find the optimal pose with the gradient descent method.
 ID-Pose can handle more than two images.
-It estimates each of the poses with additional pairs of images from triangular relationships.
-We conduct experiments using high-quality real-scanned 3D objects, where ID-Pose significantly outperforms state-of-the-art methods.
+It estimates each pose with additional pairs of images from triangular relations, where a joint optimization is applied to improve accuracy.
+We conduct extensive experiments on casually captured photos and rendered images with random viewpoints, where ID-Pose significantly outperforms state-of-the-art methods.
 Limitations.
-The running time of ID-Pose is significantly higher than neural fitting based methods.
-It requires multiple steps of pose updating by inference through large diffusion networks.
-The average time for running 100 steps is about 30 seconds on an NVIDIA V100 GPU.
-This is time-consuming and the method needs efficiency improvement.
-Due to the limitation of Zero-1-to-3, ID-Pose assumes that all cameras look at the same point and have no rotation around optical direction.
-This will introduce estimation errors for in-the-wild images.
-To extend the capability, one possible solution is to fine-tune the novel view diffusion model with the input of 6DOF poses rather than spherical coordinates.
+The execution time of ID-Pose is considerably greater than that of feed-forward networks.
+The method requires multiple iterations of pose updating through extensive diffusion networks, which is a computationally costly process.
+ID-Pose may adopt a more efficient optimization approach to improve the time efficiency.
+Due to the limitation of Zero-1-to-3, ID-Pose operates under the assumption that all cameras look at the identical point and have no rotation around the optical axis.
+While the model can accommodate minor deviations, substantial shifts may result in estimation inaccuracies.
+To mitigate this issue, a potential solution is to fine-tune the novel view diffusion model with the input of 6DOF poses rather than spherical coordinates.